OpenStudy (anonymous):

Solve the equation by finding the exact solution. ln x − ln 5 = 15

OpenStudy (anonymous):

do you know logarithm rules?

OpenStudy (anonymous):

yes because that is lnx/ln15=15?

OpenStudy (anonymous):

no.. it is supposed to be \[\ln (\frac{ x }{ 15 }) = 15\]

OpenStudy (anonymous):

oh yeah and to solve for that do e on both sides?

OpenStudy (anonymous):

yes...and what do you get?

OpenStudy (anonymous):

to have x/15=e^15?

OpenStudy (anonymous):

yes..you are right!

OpenStudy (anonymous):

what would be your final answer?

OpenStudy (anonymous):

its a huge number 49035260.59?

OpenStudy (anonymous):

you are supposed to leave your answer i the exact form.... so just leave your answer as 15e^15.

OpenStudy (anonymous):

oh i got u

OpenStudy (anonymous):

yes... so that is it!...(:

OpenStudy (anonymous):

thanks so much :)

OpenStudy (anonymous):

yw...(: